OpenStudy (anonymous):

how do you find the antiderevative of 1/sqrt(x+1)

OpenStudy (anonymous):

using substitution

OpenStudy (anonymous):

\[\int{\frac{1}{\sqrt{x+1}}}dx = \int{(x+1)^{-\frac{1}{2}}}dx\] \[u := x+1\] \[du = dx\] \[\implies \int{u^{-\frac{1}{2}}}du = 2\sqrt{u} = 2\sqrt{x+1}\]

OpenStudy (agreene):

let: u=x+1 and du=dx \[\int \frac{1}{\sqrt u}du=2\sqrt u +C= 2\sqrt {x+1}+C\]

OpenStudy (anonymous):

damn I forgot the C... the C!!!!

OpenStudy (agreene):

silly constants!

hero (hero):

I never remember the C

hero (hero):

You shouldn't have to write C. It should be implied